OpenStudy (anonymous):

solve the system of equation Re(z^2)=0,|z|=2

OpenStudy (anonymous):

\(a=a+bi\) \[Re(z^2)=0\iff a^2-b^2=0\iff a=\pm b\]

OpenStudy (anonymous):

oops i meant \[z=a+bi\]

OpenStudy (anonymous):

then \[z^2=a^2-b^2+2abi\] making \[Re(z^2)=a^2-b^2\]

OpenStudy (anonymous):

if \(a^2-b^2=0\) the \(a=b\) or \(a=-b\) so the number lives either along the line \(a=b\) or along the line \(a=-b\)

OpenStudy (anonymous):



OpenStudy (anonymous):

if \(|z|=2\) it also lives on the circle of radius \(2\)

OpenStudy (anonymous):

is it clear what those four numbers are?

OpenStudy (anonymous):

what r they?

OpenStudy (anonymous):

what would they be if it was the unit circle rather than the circle of radius 2?

OpenStudy (anonymous):

take those four answers and double them

OpenStudy (anonymous):

hint, if it was the unit circle, since one angle is clearly \(\frac{\pi}{4}\) the one in quadrant 1 would be \[\frac{\sqrt2}{2}+\frac{\sqrt2}{2}i\]

OpenStudy (loser66):

I contribute to the satellite73's stuff, that is \(\pm a=\pm b\), so that you have 4 possibilities: a =b a=-b -a = b -a=-b if |z|=2, it means |z| = \(\sqrt{a^2+b^2}\) =2 all above possibility are the same when we compute a^2 and b^2, so that just take a =b as sample a=b--> a^2+b^2 = 2a^2 --> |z| = \(\sqrt{2a^2}=2\) if and only if \(a^2 =2 --> a=\pm\sqrt{2}\) If \(a=\sqrt{2}\) and a = b, then z= a +bi = \(\sqrt{2}+\sqrt{2}i= \sqrt{2}(1+i)\)\) If \(a=\sqrt{2}\) and a =- b, then z= a +bi = \(-\sqrt{2}+\sqrt{2}i= \sqrt{2}(-1+i)\)\) do the same with the 2 possibilities left ( I mean -a=b and -a=-b), you will get the required answer